Saturday, 22 August

Saturday, 22 August

Short Communications and Poster Sessions

Short Communications: Saturday, 22 August, 14:00–14:45

14:00–14:15  14:15–14:30  14:30–14:45

Sect.: 02

Representations of Groups, Chair: Alexander S. Kleshchev

MA 311 Mohammad Reza Darafsheh The character table of the group GL_{2}(q) when extended by a certain group of order two (p.15)

Anna Helversen-Pasotto

Représentations de GL(n, F_{q}) et sommes de Gauss (p.18)

Alexei Davydov

Drinfel’d theory and representations of finite groups (p.16)

Sect.: 04

Affine Algebraic Geometry, Chair: Marko Roczen

MA 042 Tatsuji Kambayashi A Nullstellensatz for Pro-Affine Algebras

(p.55)

Vladimir Shpilrain

Embeddings of curves in the plane (p.60)

Jie-Tai Yu

Polynomial Retracts and the Jacobian Conjecture (p.63)

Sect.: 05

Surface Theory, Chair: John M. Sullivan

MA 041 Udo J. Hertrich-Jeromin Transformations of Isothermic surfaces

(p.71)

Bernd Ammann

New estimates for the Laplace and Dirac spectrum of 2-Tori and applications to the Willmore functional (p.64)

John M. Sullivan

Triunduloids: Embedded Constant-Mean-Curvature Surfaces with Three Ends

(p.81)

Sect.: 06

Algebraic Topology I, Chair: Jean-Michel Lemaire

H 112 Igor Kriz Real-oriented homotopy theory (p.91)

Carlos Prieto

Quasifibrations and Bott periodicity (p.95)

Alexander S. Mishchenko

Local combinatorial Hirzebruch formula (p.94)

Sect.: 10

Elliptic Equations III (Spectral Theory), Chair: Tetsutaro T. Shibata

H 111 Gabriella Bognár On a nonlinear eigenvalue problem (p.198)

Jochen Denzler

Existence of windows with minimal heat diffusion (p.201)

Tetsutaro T. Shibata

Asymptotic behavior of eigenvalues of nonlinear elliptic two-parameter problems

(p.221)

Sect.: 10

Systems II, Chair: Joachim Naumann

H 110 Hakop A. Hakopian On two partial differential and one algebraic systems (p.205)

Waichiro Matsumoto

On the Cauchy-Kowalevskaya theorem of Nagumo type for systems (p.213)

Joachim Naumann

Interior Hölder continuity of weak solutions to nonlinear parabolic systems in two spatial dimensions (p.215)

Sect.: 11

Field Theory I, Chair: Pierre Mathieu

H 107 Robert W. Carroll On Whitham dynamics and renormalization (p.230)

Andrei G. Bytsko

On factorized Virasoro characters (p.230)

Andreas Fring

Lie Algebraic Structures from Factorized Virasoro Characters (p.233)

Sect.: 11

Discrete Eigenvalue Theory, Chair: Miodrag Perovic

H 1029 Miodrag Perovic Some Comments Concerning the Application of the Discrete Eigenvalue Theory

Sect.: 14

Mathematical Aspects of Computer Science V, Chair: Hartmut Ehrig

MA 043 Bernd Gärtner Abstract Objective Functions — From Geometry to Combinatorics in Linear Programming (p.291)

Andreas Brieden

Optimal Radii-Algorithms (p.289)

Anuar Dusembaev

An operator approach to program segmentation (p.290)

Sect.: 15

Partial Differential Equations IV, Chair: Jennifer J. Zhao

H 2033 Jennifer J. Zhao Finite-difference approximations to a parabolic system modeling the thermoelastic contacts of two rods

(p.319)

Jianping Zhu

A High Order Algorithm for Solving Reaction Diffusion Equations (p.319)

Raul Kangro

Analysis of Artificial Boundary Conditions for Black-Scholes Equations (p.304)

Sect.: 17

Optimization in Function Spaces, Chair: Valeri V. Obukhovskii

H 2013 Andres, M. Encinas Existence and Approximation of Gauss Variational Problem in Potential Theory

(p.340)

Stanisław Migórski

Optimal control of hemivariational inequalities (p.345)

Valeri V. Obukhovskii

On some methods in a periodic problem for a functional differential inclusion in a Banach space (p.346)

Short Communications: Saturday, 22 August, 15:00–15:45

15:00–15:15  15:15–15:30  15:30–15:45

Sect.: 02

Commutative Rings, Chair: Peter Draexler

MA 311 Alexander Levin Multivariable dimension polynomials and their invariants (p.21)

Venkat Murali

On Krull Intersection Theorem of Fuzzy Ideals (p.24)

Cornelia G. Naude

Asymptotic behaviour of attached primes (p.25)

Sect.: 04

Riemann Surfaces/Linear Systems, Chair: Anette A’Campo-Neuen

MA 042 Sergei Natanzon Topological structure of spaces of holomorphic morphisms (p.58)

Masahiro Ohno

On nef values of determinants of ample vector bundles (p.59)

Siye Wu

On the instanton complex of holomorphic Morse theory (p.62)

Sect.: 05

Curves, Chair: Joachim Giesen

MA 041 Kazuyuki Enomoto Total absolute curvature of nonclosed curves (p.67)

Joachim Giesen

Geometry of One-Manifolds (p.69)

Sect.: 06

Algebraic Topology II, Chair: Daniel H. Gottlieb

H 112 Carles Casacuberta A universal space for plus-constructions

(p.86)

Samson Saneblidze

Adams Cobar Construction as a Cubical Complex (p.97)

Hirotaka Tamanoi

BP cohomology theory and v_{n}-series (p.98)

Sect.: 10

Hyperbolic Equations II (Shocks, Blow-up, and Other Singular Phenomena), Chair: Ya-Guang Wang

H 111 Mehmet Can Blow up of solutions of a quasilinear hyperbolic equation with applications (p.199)

Michael Ruzicka

Electrorheological fluids, an example for systems with non-standard growth (p.219)

Ya-Guang Wang

Nonlinear Geometrical Optics for Shock Waves (p.224)

Sect.: 10

Diffusion Processes, Chair: Natalia I. Grinberg

H 110 Jack Carr Asymptotic Behaviour of Solutions to a simplified Lifshitz-Slyozov Equation

(p.200)

Alexander M. Chebotarev

Resolvent theory for quantum dynamical semigroup (p.200)

Natalia I. Grinberg

Inverse boundary problem for 2D diffusion equation (p.205)

Sect.: 11

Field Theory II, Chair: Robert W. Carroll

H 107 Pierre Mathieu Generating functions for fusion rules

(p.240)

Moukadas Missarov

Critical phenomena and continuum limit in the hierarchical fermionic model (p.240)

Michael Monastyrsky

Correlation functions on Riemann surfaces and statistics of Knots (p.241)

Sect.: 14

Mathematical Aspects of Computer Science VI, Chair: Andreas Brieden

MA 043 Hartmut Ehrig On the Role of Category Theory in Computer Science (p.290)

Raja Natarajan

Combinators for Concurrency (p.294)

George Francis

The Making of the Optiverse (p.291)

Sect.: 15

Partial Differential Equations V, Chair: Cesar Palencia

H 2033 Cesareo González The implicit Euler method for abstract time-dependent parabolic problems with variable domains (p.301)

Raimondas Ciegis

Odd-even difference schemes for the Schrödinger problem (p.298)

Daisuke Furihata

Finite difference schemes that inherit the energy conservation or dissipation property from partial differential equations (p.300)

Sect.: 17

Stochastic Control/Controllability, Chair: Galina A. Timofeeva

H 2013 Yuri L. Sachkov Controllability of right-invariant systems on solvable Lie groups (p.349)

Leonid E. Shaikhet

Stability of stochastic difference equations with varying delays (p.350)

Galina A. Timofeeva

Effective control and observation in the multistage stochastic optimization problem (p.352)

Short Communications: Saturday, 22 August, 16:15–17:00

16:15–16:30  16:30–16:45  16:45–17:00

Sect.: 03

Integral Matrices, Chair: Jean-Louis Nicolas

MA 311 Shashikant Katre Matrices over orders in algebraic number fields as sums of k-th powers (p.39)

Dmitri Malinin

Integral representations of finite groups (p.42)

Jean-Louis Nicolas

On the maximum order of torsion elements in GL(n, ℤ) (p.43)

Sect.: 03

Algebraic Geometry II, Chair: Jörg Wildeshaus

MA 042 Jörg Wildeshaus Degeneration of Shimura Polylogarithms

Sect.: 07

Topological Groups, Chair: Wolfgang Reinhold Bergmann

H 112 Boris S. Khots Fifth Hilbert problem and infinitedimensional P-groups (p.106)

Grigori Lazarevich Litvinov

Group algebras and nonunitary representations of smooth groups (p.107)

Julio Cesar Rebelo

Associating vector fields to certain subgroups of Diff^{w}(S^{1}): applications to ergodicity and rigidity (p.109)

Sect.: 08

Univalent Functions II, Chair: Arne Stray

H 110 Peter Duren Hyperbolic Capacity and its Distortion under Conformal Mapping (p.122)

Ilgiz G. Kayumov

Asymptotical behaviour of logarithmic coefficients of univalent functions (p.134)

Dmitri V. Prokhorov

Estimates of linear functionals for bounded univalent functions (p.149)

Sect.: 08

Inequalities and Interpolation, Chair: Vladimir I. Ovchinnikov

MA 041 Vladislav F. Babenko Inequalities of Landau–Kolmogorov–Nagy Type (p.115)

Vladimir Alexandrovich Kofanov

Exact Inequalities of Kolmogorov Type for Periodic Functions of Small Smoothness

(p.135)

Vladimir I. Ovchinnikov

Nonclassical Interpolation Theorems in Spaces of Smooth Functions (p.146)

Sect.: 08

Measure and Integration, Chair: Carmen D. Vlad

MA 309 Alexandre V. Mikhalev Solution of Riesz-Radon problem (p.141)

Valeriy Konstantinovich Zakharov

A solution of the general Riesz-Radon problem for a Hausdorff space (p.166)

Carmen D. Vlad

Outer measures and countably additivity of measures (p.162)

Sect.: 08

Banach Algebras and Fredholm Maps, Chair: Kari E. Ylinen

MA 043 Victor G. Zvyagin Topological methods in the theory of perturbations of Fredholm maps and applications (p.167)

Hans-Olav I. Tylli

Rigidity of commutators and elementary operators on Calkin algebras (p.161)

Kari E. Ylinen

Positive definite operator functions on groups and related multilinear operators on products of Banach-^{∗}-algebras (p.165)

Sect.: 09

Random Dynamical Systems, Chair: Govindan Rangarajan

H 111 Lev Ryashko Quasipotential in stochastic stability analysis of limit cycles (p.188)

Henri U. Schurz

Contractivity and Stability Exponents of Nonlinear Stochastic Dynamical Systems

(p.191)

Govindan Rangarajan

A New Method for Computing Lyapunov Exponents (p.186)

Sect.: 12

Self Similarity, White Noise, Chair: Murad S. Taqqu

H 107 Iosif Kolumban Invariant sets in Menger spaces (p.257)

Louis F. Rossi

Self-similar viscous rotating vortical structures (p.268)

Yuh-Jia Lee

Generalized White Noise Functionals on Classical Wiener Space (p.258)

Sect.: 12

Tests and Confidence Sets, Chair: Ekkehard Glimm

H 1029 Alejandra Cabana Goodness-of-Fit Tests Based on Transformed Empirical Processes (p.249)

Roman Rozanski

Asymptotically consistent and efficient fixed-size sequential confidence sets

(p.269)

Ekkehard Glimm

Power and admissibility of spherical multivariate tests (p.251)

Sect.: 13

Geometry III, Chair: Klaus Altmann

H 106 Jennifer Mc Nulty Connected Hyperplanes in Binary Matroids

(p.282)

Andreas W.M. Dress

Highly Regular Spatial Structures: Theory, Software, and Visualization (p.276)

Klaus Altmann

Hodge numbers for polytopes (p.274)

Sect.: 18

Undergraduate Mathematics at the University, Chair: Joseph W. Eyles

H 2013 Talma O. Leviatan Misconceptions, fallacies, pitfalls, and paradoxes in the teaching of probability

(p.361)

Lyudmila Y. Khilchenko

An algorithmic approach to the notion of an elementary function (p.360)

Joseph W. Eyles

Discovering R.L. Moore’s Calculus Course (p.358)

Short Communications: Saturday, 22 August, 17:15–18:00

17:15–17:30  17:30–17:45  17:45–18:00

Sect.: 03

Analytic Number Theory I, Chair: Jin Nakagawa

MA 311 Antanas Laurincikas Value-distribution theorems on Lerch zetafunctions (p.40)

Roberto J. Miatello

Estimates for Kloosterman sums for totally real number fields (p.42)

Jin Nakagawa

Ohno conjecture on the zeta functions associated with the space of binary cubic forms (p.43)

Sect.: 04

Curves and Surfaces/Higher-Dimensional Varieties I, Chair: Rolf-Peter Holzapfel

MA 042 Viktor S. Kulikov On Chisini’s Conjecture (p.55)

Kazuhisa Maehara

Deformation of function fields (p.56)

Dan Avritzer

The Hilbert scheme component of the intersection of two quadrics (p.50)

Sect.: 07

Amenable Groups, Chair: Grigori L. Litvinov

H 112 Wolfgang Reinhold Bergmann Amenable Groups and Generalized Cuntz Algebras (p.102)

Alexander I. Shtern [Chtern]

Quasi-Symmetry and Amenability (p.110)

Sect.: 08

Hilbert and Krein Spaces, Chair: Vladimir Strauss

H 110 Elias Flytzanis An invariance property of quadratic forms

(p.123)

Michael Gebel

Spectral Theory of Adjointable Operators in Spaces with two Metrics (p.125)

Vladimir Strauss

On a model representation for a normal operator in Krein space (p.158)

Sect.: 08

Multivalued Mappings, Weighted Inequalities, and Means, Chair: Lars-Erik Persson

MA 041 B.K. Sharma A Strong Convergence Theorem for Multivalued Nonexpansive Mappings (p.155)

Zheng Liu

Inequalities for Some Means (p.139)

Lars-Erik Persson

Sharp multidimensional multiplicative inequalities for weighted Lebegue spaces

(p.147)

Sect.: 08

Dunkl Operators and Hyperfunctions, Chair: Dohan Kim

MA 309 Margit Rösler Positivity of Dunkl’s intertwining operator

(p.152)

Stevan Pilipović

Suppleness of Some Quotient Spaces of Microfunctions (p.147)

Dohan Kim

Moment problem for hyperfunctions and the heat kernel method (p.134)

Sect.: 08

Measure Problems, Chair: Jiří Jarník

MA 043 Raymond J. Grinnell Two real functions associated with convolutions of measures (p.127)

Daniar H. Mushtari

Orthomodular posets of projections on linear spaces and signed measures on them

(p.143)

Jiří Jarník

Convergence theorem for Henstock– Kurzweil integral based on the convergence of primitives (p.132)

Sect.: 09

Complex Differential Equations, Chair: Bruno Scardua

H 111 Ljudmila A. Bordag Projective differential geometrical structure of the Painlevé equations (p.172)

Alexei A. Glutsuk

Stokes phenomena via generic perturbation (p.176)

Bruno Scardua

Integration of Complex Differential Equations (p.190)

Sect.: 12

Teaching in Probability, Chair: Murad S. Taqqu

H 107 Natalya V. Vinogradova Classical Definition of Probability (p.271)

Yuliya Stepanovna Mishura

Generalized stochastic integrals with application to fractional Brownian motion

Sect.: 13

Codes and Cryptography, Chair: A. Donald Keedwell

H 106 Gerhard R. Paseman A different approach to Hadamard’s maximum determinant problem (p.283)

Gholamrega B. Khosronshahi

On trades and their codes (p.279)

A. Donald Keedwell

Crossed-inverse quasigroups and their use in cryptography (p.278)

Sect.: 17

Nonlinear Optimization II, Chair: Leon A. Petrosjan

H 2033 Alexei V. Blinov The combined methods of centers with using reduced gradient for solving nonlinear optimization problems (p.337)

R. Enkhbat

Global Optimality Conditions for some class of Nonconvex Optimization Problems

(p.340)

Leon A. Petrosjan

Time-consistancy problem in differential games (p.348)

Sect.: 18

Geometry in the Classroom, Chair: F. Rudolf Fritsch

H 2013 Klaus-D. Graf Distance learning fördert die Kreativität im Mathematikunterricht (p.360)

Bernd Zimmermann

Al Sijzi on problemsolving and creativity in elementary geometry (p.364)

F. Rudolf Fritsch

Remarks on quadrangles (p.359)

Poster Session: Saturday, 22 August, 14:00–15:45

Sect.: 12

Probability and Statistics

MA 144 Jaap Geluk Second order regular variation and the domain of attraction (p.251)

Werner S. Hurlimann

Comparison between extremal distributions: a conjecture and a partial proof

(p.253)

Takeshi Ishikawa

On modified maximum likelihood estimation for Rayleigh distribution (p.254)

Dionis Lica χ^{2}-Test for Markov Chains (p.258)

Alfred Liemant

Hopping transport of charge carriers on a disordered lattice (p.259)

Elina M. Moldavskaya

Minimum Contrast Estimators of Parameter of Spectral Density of Random Fields in Continuous Case (p.261)

Hans-Joachim Mucha Classification based on adaptive distances

(p.262)

Marcus L.G. Nilsson

Finite Probability Sequences of Large Entropy (p.264)

Jörg Polzehl

Adaptive Image Denoising with Applications to MRI (p.266)

Tiberiu V. Postelnicu Specific Statistical Method for Binary Data

(p.266)

Grzegorz A. Rempala

Random permanents of increasing degree (p.267)

Mikhailo Yadrenko

Extrapolation and filtering problems for isotropic random fields (p.273)

Sect.: 18

Teaching and Popularization of Mathematics

MA 144 Gloria L.M. Alves Modernization of Basic Disciplines in Engineering Courses at IME (p.357)

Gerhard Brecht

Do Mathematical Standards Concern Mathematics? (p.357)

Cecilia Crespo Crespo

Thinking about Geometry Teaching (p.358)

Paul Fjelstad A Picture for Real Arithmetic (p.359)

Peter Kortesi

Junior Mathematical Congress — Challenge for Future Mathematicians (p.360)

Elena Vladimirovna Kuragina

Methodical support for a programmingmethodical complex of study methods of optimization (p.361)

Julia M. Moshnikova Correspondence School of Mathematics in Karelia (p.362)

Yasuhiro Ozaki

An Approach in Teaching by Means of the Personal Computer Software “Mathematica” (p.363)

Poster Session: Saturday, 22 August, 16:15–18:00

Sect.: 14

Mathematical Aspects of Computer Science

MA 144 Giovanni Francesco Mascari From TQFT to TQCT (p.293)

Roman Matuszewski

Journal of Formalized Mathematics — a collection of mechanically checked articles

(p.293)

Sect.: 15

Numerical Analysis and Scientific Computing

MA 144 Mansour Abdullah Al-Zanaidi A Class of Generalized Trapezoidal Formulas for the Integration of y’=f(x, y) (p.296)

Mehmet Bozer

The composite grids method of increased accuracy in solving the mixed boundary value problem for the Laplace equation

(p.297)

Chin-Yun Chen

VANI: A package for Validated, Adaptive Numerical Integration in one and more dimensions (p.297)

Adigezal Dosiyev A block-grid method for solving Laplace’s boundary value problem with singularities

(p.298)

Bratislav Iricanin

On the modification of Euler-Maclaurin summation formula connected with the Chebyshev quadrature formulae (p.302)

Anna Kamont

Ditzian — Totik moduli of smoothness, spline systems and generalized Hölder spaces (p.303)

Ivan G. Kazantsev Tomographic reconstruction from arbitrary directions (p.304)

Ingemar Kinnmark

Overdetermined Differential Equation Problems (p.304)

Peeter Oja

Adaptive by monotonicity interpolating splines (p.307)

Arvet Pedas Collocation method for weakly singular integral equations (p.309)

Gerd Reinhardt

Schnelle Algorithmen zum Clustern großer Datenmengen (p.311)

Mayumi Shoji

Bifurcation structure of steady, deepwater gravity waves with constant vorticity

(p.313)

Svetlana N. Sytova Mathematical Modelling of X-ray Dynamical Diffraction by Non-Stationary Deformed Crystals (p.315)

Zoltan Istvan Szabo

Root-finding Inclusion Algorithms Based on Tangential Lines and Parabolas (p.315)

Svetlana B. Zaitseva

On sharp error estimates in L_{2}-norm for methods of solving the heat equation in non-smooth data classes (p.318)

Aihui Zhou Multiparameter error resolution for finite elements (p.319)